Use the following equation:
3 C2H3O2H + Al(OH)3 + Al(C2H302)3 + 3 H2O
How many moles of H2O can be made with 1.59 mol of Al(OH)3?
1
Expert's answer
2020-03-02T06:56:32-0500

n(H2O) = n(Al(OH)3) * 3 = 1.59 * 3 = 4.77 mol